# Latency Adjustable Transformer Encoder for Language Understanding

### v8

@@ III-D Proposed Architecture
According to the discussion in Section III-B, almost all latency due to the FLOPs occur in the encoder layers. In addition, according to the observation in Section III-C, the word-vector contribution in the encoder decreases at later layers. Based on these observations, this Section proposes the details of the proposed architecture.
The proposed method is applied to the BERTbase model for explanation and evaluation. The proposed architecture is the same as the original BERT with a modified multi-head self-attention layer. Fig. 5 illustrates the proposed multi-head self-attention layer. The proposed architecture adds the Sort and Eliminate layers in the middle of the original BERT multi-head self-attention layer. The Sort layer sorts the input context according to the computed layer Score Vector. The Eliminate layer eliminates a fraction of the sorted context with lower scores according to the Elimination-Rate ($\alpha_{ER}$).
@@
The effective number of processed word-vectors can be estimated using the number of word-vectors before and after the eliminate layer. Based on Table I, in each encoder layer, around 75% of FLOPs occur after the self-attention layer. Equation (4) presents the effective number of Processed word-vectors ($PW_{l}$) in layer $l$.
$PW_{l}\approx\frac{T_{l-1}+3T_{l}}{4}$ (4)
@@
$PW\approx T\left[\frac{1}{4}+\sum_{i=1}^{L-1}\prod_{j=1}^{i}\alpha_{EP}^{j}+\frac{3}{4}\prod_{j=1}^{L}\alpha_{EP}^{j}\right]$ (5)
Processed word-vectors in the original BERTbase is:
[…]
In these equations, ${L}$ is the number of BERT encoder layers, and ${T}$ is the number of encoder input word-vectors. Since almost all inference latency occurs in the encoder, the overall model computation effort is almost equal to that of the encoder stage. Equation (7) presents the proposed method inference time speedup to BERTbase.
$K_{speedup}\approx\frac{4L}{1+4\sum_{i=1}^{L-1}\prod_{j=1}^{i}\alpha_{EP}^{j}+3\prod_{j=1}^{L}\alpha_{EP}^{j}}$ (7)
The above equation assumes that the number of remaining word-vectors in each encoder layer is more than a single word-vector. In other words, (3) returns the second argument, which is valid for most cases. For example, if $L$ is 12 and $\alpha_{EP}^{l}$ is constant and equal to 0.8 for all layers, the method inference speed is around 3.09 times faster than BERTbase.
@@
III-E Extending Proposed Architecture to Decoder Models
The applicability of the proposed method can be extended to non-encoder Transformers, such as the GPT-family Large Language Models (LLMs). GPT-2 [71] represents a decoder-based transformed language model widely utilized in text generation tasks. This section outlines the considerations when implementing the proposed method on the GPT-2 model.
The foundational GPT-2 model comprises 12 decoder layers, with each decoder featuring a causal attention layer equipped with 12 heads. The total number of parameters in the GPT-2 model is approximately 117 million. A key distinction between the architectures of GPT-2 and the BERTbase model lies in the design of their attention layers. As explained in Section III-A, the attention layers of BERTbase encoders are bidirectional, involving all input word-vectors in generating each output word-vector. Conversely, the GPT-2 model, tailored for text generation tasks, employs causal attention layers, where only the preceding word-vectors contribute to the attention layer output at each time stamp.
In a causal attention layer, the attention probability matrix, represented as $M_{Att}^{l}$, adopts the form of a triangular matrix. To extend the proposed architecture outlined in Section III-D to the GPT-2 model, an additional normalization step needs to be integrated after the fifth step in Algorithm 1. This supplementary step normalizes the Score Vector $M_{S}^{l}$ based on the count of non-zero elements in the probability matrix $M_{Att}^{l}$. Equation (8) articulates this normalization process in layer $l$, wherein the $i$-th element of the Score Vector is adjusted relative to the number of non-zero elements in the $i$-th column of the probability matrix. Through this normalization step, the resulting score values become comparable.
@@ IV Experiments
In this section, we implement the proposed method on both BERTbase, GPT-2 and Flan-T5 [72] models, delving into the method’s effectiveness across various natural language understanding tasks, text generation datasets and instruction tuning benchmarks. Towards the conclusion of this section, we delve into the offline-tuning property and conduct an ablation study on the proposed self-attention layer.
@@ IV-A Datasets
The proposed method performance is assessed in encoder models using the IMDB [62] dataset, which is a sentiment analysis task, as well as the General Language Understanding Evaluation (GLUE) [63] benchmark, which consists of two single-sentence classification tasks: SST-2 [64] and CoLA [73], two multi-sentence similarity classification tasks: MRPC [65] and QQP [66], one binary question answering task: QNLI [67], three natural language inference tasks: RTE [68], MNLI-M and MNLI-MM [69], and one multi-sentence similarity regression task: STS-B [70].
To assess the performance of the method in decoder models, we conduct evaluations on several datasets. The WikiText-103 [74] language modeling dataset, comprising over 100 million tokens extracted from verified Good and Featured articles on Wikipedia, serves as a primary evaluation ground. Additionally, the PTB [75] dataset, featuring a million words from 1989 Wall Street Journal material, and the 1BW [76] dataset, established as a benchmark corpus for measuring progress in statistical language modeling, are included in our evaluation. Moreover, we assess the method on The LAMBADA [77] dataset, a collection of narrative passages characterized by the intriguing feature that human subjects can predict their last word when exposed to the entire passage. Extracted from BookCorpus, The LAMBADA dataset comprises 10,022 passages, contributing to a comprehensive evaluation of the proposed method’s performance in decoder models.
To more closely show the method’s performance to understand language, it is applied to the Flan-T5 model and evaluated on the Massive Multitask Language Understanding (MMLU) benchmark, which includes exam questions from 57 tasks such as mathematics, history, law, and medicine [78], and the Big Bench Hard (BBH) benchmark, which includes 23 challenging tasks [79].
IV-B Model Settings
@@
For the instruction tuning evaluation, the Flan-T5 model is used. Flan-T5 is an enhanced version of T5 [80] that has been fine-tuned on a mixture of tasks. The Flan-T5 encoder self-attention layers are substituted with the proposed encoder self-attention layer, and other layers are modified to work with a dynamic number of word vectors in each encoder and decoder layer. The Flan-T5 weights are directly used in our model without any fine-tuning process. the $\alpha_{ER}$ is selected in the range [0.8, 1.0], and other evaluation hyper-parameters are the same as those of the default Flan-T5 model.
@@ IV-C Implementation Details
The initial models utilized for this study are the pre-trained versions of BERTbase-uncased [4], GPT-2 [71], and Flan-T5 [72] from the Hugging Face [^2] https://huggingface.co/ implementations. The implementation is carried out in the PyTorch [^3] https://pytorch.org/ library, utilizing an Nvidia Tesla P100 GPU for BERTbase, an Nvidia Tesla T4 GPU for GPT-2, and an Nvidia Tesla A100 GPU for Flan-T5. The datasets are sourced from the Hugging Face Datasets library, and the GLUE results are validated through the official GLUE service [^4] https://gluebenchmark.com/. The instruction tuning assessment is done with help of the INSTRUCTEVAL [81] method. This framework ensures consistency and reliability in the assessment of the proposed method’s performance on the specified models.
IV-D Experimental Results of BERTbase
The experimental results for extensive tasks are presented in Table II. This work is 2.9 times faster than BERTbase, with 0.7% drop on the test set. Also, the model inference time is up to 4.8 times faster than BERTbase, with only 0.9% accuracy drop. These results indicate that the proposed method effectively improves Transformer-based models’ performance and resource usage with minimal accuracy degradation.
Table II shows the experimental speedup, the expected speedup, and the number of FLOPs for each task. The number of FLOPs reduction is consistent with inference speedup and expected speedup predicted using (7). The slight differences are due to the computational overheads, including the computational effort of the tokenizer and output classifier layers that are not considered in (7). The consistency between the experimental speedup, the expected speedup, and the FLOPs reduction approves the hypothesis discussed in Section III-D.
TABLE II: The Proposed Method Experimental Results on GLUE [63] benchmark and IMDB [62] dataset
Model CoLA SST-2 MRPC QQP QNLI RTE MNLI-M MNLI-MM STS-B IMDB Avg
Test Resulta BERTbase 52.7 93.5 88.7 71.2 91.0 68.9 84.7 83.8 84.3 93.7 81.25
@@ IV-D Experimental Results of BERTbase
Inference Speedup - 3.8 2.4 2.9 4.7 2.3 2.6 2.5 2.5 2.3 2.7 2.9
Expected Speedupb - 3.67 2.35 2.88 4.76 2.25 2.64 2.49 2.49 2.22 2.63 2.84
#FLOPs - 5.7B 9.1B 7.6B 4.7B 9.8B 8.5B 8.7B 8.7B 9.5B 8.6B 7.7B
Input Seq Length - 64 64 128 128 128 256 128 128 64 512 160
a Matthew’s Correlation is reported for CoLA; F1-score for MRPC and QQP; Pearson Correlation for STS-B, and Accuracy for the rest.
b With using equation (7).
TABLE III: Results comparison between BERTbase [4], DistilBERT [46], TinyBERT [47], MobileBERT [48], PoWER-BERT [25], LayerDrop [39], FPT [34] and BERTbase with latency adjustable encoder (This Work)
Modela Method CoLA SST-2 MRPC QQP QNLI RTE MNLI-M MNLI-MM Avg Drop FLOPs Speedup
BERTbase - 52.7 93.5 88.7 71.2 91.0 68.9 84.7 83.8 0.00 22.5B 1
DistilBERT KD 51.3 91.3 87.5 70.1 89.2 59.9 82.2 - 2.74 11.3B 1.7
TinyBERT KD 51.1 93.1 87.3 71.6 90.4 70 84.6 83.2 0.40 11.3B 2.0
MobileBERTb KD 50.5 92.8 88.8 70.2 90.6 66.2 83.3 82.6 1.19 5.7B 5.5
PoWER-BERTc Pruning 52.3 92.1 88.1 70.2 90.1 67.4 83.8 83.1 0.92 - 2.9
LayerDrop Pruning - 92.5 - - 89.4 - 82.9 - 1.47 11.3 1.7
FPT Pruning - 92.6 - - 90.3 - 83.5 - 0.93 14.5 1.5
This Work Pruning 53.4 92.6 88.0 70.1 90.2 67.9 83.7 82.7 0.73 7.7B 2.9
a Matthew’s Correlation reported for CoLA; F1-score for MRPC and QQP; Pearson Correlation for STS-B; Accuracy for the rest.
b MobileBERT starting point is from BERTlarge architecture that is around three times more parameters with respect to BERTbase.
c PoWER-BERT training consists of 3 fine-tuning steps, and the model inference speedup is fixed at inference.
Table III compares the proposed method and the previously reported methods. BERTbase model ($L$=12, $h$=12, $d$=768) is used as the basis for comparison. The test results, the average accuracy drop compared to BERTbase, the average number of FLOPs, and the average inference speedup are reported for each model. As shown in the table, the proposed method achieves 2.9 times speedup with 0.73% on average accuracy drops compared to BERTbase model.
@@
The table shows that the results of the proposed method are competitive and more effective than previous baseline methods across a wide range of experimental tasks. Furthermore, compared to recently reported methods, our approach is applied only during the fine-tuning phase and offers the offline-tuning property. Blank fields in the table, indicate instances where some works did not report their assessments on certain datasets.
IV-E Experimental Results of GPT-2
To assess the impact of the proposed method on the overall context of the input, the approach is implemented on the GPT-2 model following the procedures detailed in Section III-E. The training process involves utilizing 1% of the OpenWebText [82] dataset with a stride of 128 for three epochs. This methodology is employed to comprehensively evaluate how the proposed method influences the global context of the input when applied to the GPT-2 model.
Table IV presents the results of the zero-shot perplexity assessment for the proposed method in text generation tasks, offering a basis for comparison with other contemporary studies. The findings demonstrate that the proposed method effectively maintains the global context of the input sequence. The proposed Attention Context Contribution (ACC) metric performs well, ensuring that the elimination of word-vectors does not compromise the input’s global context. It is noteworthy that the original GPT-2 model, trained by OpenAI [71], benefited from more extensive hardware and dataset resources. However, our proposed architecture enables efficient adaptation with just a few small epochs. Furthermore, the proposed method exhibits superior performance when compared to the DistilGPT2 [46] and ZipGPT2 [83] models in the evaluation.
TABLE IV: Perplexity (PPL) comparison between GPT-2 [71], DistilGPT2 [46] and ZipGPT2 [83] with latency adjustable GPT-2

### v9

@@ III-D Proposed Architecture
According to the discussion in Section III-B, attention and feed-forward layers are major contributes of the total latency. In addition, according to the observation in Section III-C, the word-vector contribution in the encoder decreases at later layers. Based on these observations, this Section proposes the details of the proposed architecture.
In proposed architecture, the orginal BERT self-attention layer is modified with suggested self-attention layer. Fig. 5 illustrates the proposed multi-head self-attention layer. The proposed architecture adds the Sort and Eliminate layers in the middle of the original BERT multi-head self-attention layer. The Sort layer sorts the input context according to the computed layer Score Vector. The Eliminate layer eliminates a fraction of the sorted context with lower scores according to the Elimination-Rate ($\alpha_{ER}$).
@@
The effective number of processed word-vectors can be estimated using the number of word-vectors before and after the eliminate layer. Based on Table I, in BERT-base, around 65% of latency occur after the self-attention layer. Equation (4) presents the effective number of Processed word-vectors ($PW_{l}$) in layer $l$.
$PW_{l}\approx\frac{7T_{l-1}+13T_{l}}{20}$ (4)
@@
$PW\approx T\left[\frac{7}{20}+\sum_{i=1}^{L-1}\prod_{j=1}^{i}\alpha_{EP}^{j}+\frac{13}{20}\prod_{j=1}^{L}\alpha_{EP}^{j}\right]$ (5)
Processed word-vectors in the original BERT-base is:
$PW_{BERT-base}=T\times L$ (6)
In these equations, ${L}$ is the number of BERT encoder layers, and ${T}$ is the number of encoder input word-vectors. Since almost all inference latency occurs in the encoder, the overall model computation effort is almost equal to that of the encoder stage. Equation (7) presents the proposed method inference time speedup to BERT-base.
$K_{speedup}\approx\frac{20L}{7+20\sum_{i=1}^{L-1}\prod_{j=1}^{i}\alpha_{EP}^{j}+13\prod_{j=1}^{L}\alpha_{EP}^{j}}$ (7)
The above equation assumes that the number of remaining word-vectors in each encoder layer is more than a single word-vector. In other words, (3) returns the second argument, which is valid for most cases. For example, if $L$ is 12 and $\alpha_{EP}^{l}$ is constant and equal to 0.8 for all layers, the final inference speed is around 2.99 times faster than BERT-base.
@@
III-E Extending Proposed Architecture to Autoregressive Models
The suggested solution can also be applied to decoder architectures, including GPT, Gemma, Mistral, and Llama Large Language Models (LLMs). This section outlines the key considerations for implementing the method on the GPT-2 model, as a representative example of autoregressive models.
The foundational GPT-2 model comprises 12 decoder layers, with each decoder featuring a causal attention layer equipped with 12 heads. The total number of parameters in the GPT-2 model is approximately 124 million. A key distinction between the architectures of GPT-2 and the BERT-base model lies in the design of their attention layers. As explained in Section III-A, the attention layers of BERT-base encoders are bidirectional, involving all input word-vectors in generating each output word-vector. Conversely, the GPT-2 model, tailored for text generation tasks, employs causal attention layers, where only the preceding word-vectors contribute to the attention layer output at each time stamp.
In a causal attention layer, the attention probability matrix, represented as $M_{Att}^{l}$, adopts the form of a triangular matrix. To extend the proposed architecture outlined in Section III-D to causal attention layers, an additional normalization step needs to be integrated after the fifth step in Algorithm 1. This supplementary step normalizes the Score Vector $M_{S}^{l}$ based on the count of non-zero elements in the probability matrix $M_{Att}^{l}$. Equation (8) articulates this normalization process in layer $l$, wherein the $i$-th element of the Score Vector is adjusted relative to the number of non-zero elements in the $i$-th column of the probability matrix. Through this normalization step, the resulting score values become comparable.
@@ IV Experiments
In this section, we implement and evaluate the suggested method on BERT-base, GPT-2, Flan-T5 [86], Gemma2 [75], Mistral, and Llama3 [76], exploring its effectiveness across various natural language understanding tasks, text generation datasets, and instruction-tuning benchmarks. Toward the conclusion of this section, we examine the offline-tuning property and conduct an ablation study on the proposed attention layer.
@@ IV-A Datasets
The performance of the proposed method on natural language understanding tasks is evaluated using the IMDB dataset [77], which focuses on sentiment analysis, as well as the General Language Understanding Evaluation (GLUE) benchmark [78]. The GLUE benchmark includes two single-sentence classification tasks: SST-2 [79] and CoLA [87]; two multi-sentence similarity classification tasks: MRPC [80] and QQP [81]; one binary question answering task: QNLI [82]; three natural language inference tasks: RTE [83], MNLI-M, and MNLI-MM [84]; and one multi-sentence similarity regression task: STS-B [85].
To assess the performance of the approach on text generation tasks, we conduct evaluations on several datasets. The WikiText-103 [88] language modeling dataset, comprising over 100 million tokens extracted from verified Good and Featured articles on Wikipedia, serves as a primary evaluation ground. Additionally, the PTB [89] dataset, featuring a million words from 1989 Wall Street Journal material, and the 1BW [90] dataset, established as a benchmark corpus for measuring progress in statistical language modeling, are included in our evaluation. Moreover, we assess the approach on The LAMBADA [91] dataset, a collection of narrative passages characterized by the intriguing feature that human subjects can predict their last word when exposed to the entire passage. Extracted from BookCorpus, The LAMBADA dataset comprises 10,022 passages, contributing to a comprehensive evaluation of the proposed method’s performance in decoder models.
To further demonstrate the suggested approach’s performance in language understanding, it is evaluated on instruction-tuning tasks and assessed using the Massive Multitask Language Understanding (MMLU) benchmark. This benchmark includes exam questions from 57 diverse tasks, covering subjects such as mathematics, history, law, and medicine [92].
IV-B Models Settings
@@
For the instruction-tuning evaluation, the Flan-T5111Flan-T5 is an enhanced version of T5 [74], Gemma2, Mistral, and Llama3 models are targeted. In these models, the self-attention layers are replaced with the proposed self-attention layer, and other layers are adjusted to accommodate a dynamic number of word vectors in each encoder and decoder layer. The models use their default weights without any tuning or fine-tuning. The parameter $\alpha_{ER}$ is selected within the range [0.8, 1.0], and all other evaluation hyperparameters remain consistent with those of the default models.
@@ IV-C Implementation Details
The initial models utilized for this study are the pre-trained versions of BERT-base, GPT-2, Flan-T5, Gemma2, Mistral, and Llama3 from the Hugging Face [^2] https://huggingface.co/ implementations. The implementation is carried out in the PyTorch [^3] https://pytorch.org/ library, utilizing an Nvidia Tesla P100 GPU for BERT-base, an Nvidia Tesla T4 GPU for GPT-2, and an Nvidia Tesla A100 GPU for Flan-T5, Gemma2, Mistral, and Lllama. The datasets are sourced from the Hugging Face Datasets library, and the GLUE results are validated through the official GLUE service [^4] https://gluebenchmark.com/. The instruction tuning assessment is done with help of the INSTRUCTEVAL [93] method. This framework ensures consistency and reliability in the assessment of the proposed solution’s performance on the specified models.
IV-D Experimental Results on Language Understanding Tasks
The experimental results of the proposed method applied to BERT-base for a wide range of tasks are presented in Table II. The proposed approach is 2.9 times faster than BERT-base, with only a 0.7% accuracy drop on the test set. Additionally, the model’s inference time is up to 4.8 times faster than BERT-base, with just a 0.9% reduction in accuracy. These results demonstrate that the suggested method significantly enhances the performance and resource efficiency of Transformer-based models while maintaining minimal accuracy loss.
The table shows the experimental speedup, and the expected speedup. The inference speedup is consistent with expected speedup predicted using (7). The slight differences are due to the computational overheads, including the computational effort of the tokenizer and output classifier layers that are not considered in (7). The consistency between the experimental speedup, the expected speedup, approves the hypothesis discussed in Section III-D.
TABLE II: Experimental Results of the Proposed Method on the GLUE [78] Benchmark and IMDB [77]
Dataset with BERT-base as the Target Model
Dataseta CoLA SST-2 MRPC QQP QNLI RTE MNLIM MNLIMM STS-B IMDB Avg
BERT-base 52.7 93.5 88.7 71.2 91.0 68.9 84.7 83.8 84.3 93.7 81.25
@@ IV-D Experimental Results on Language Understanding Tasks
Expected Speedupb 3.7x 2.4x 2.9x 4.8x 2.3x 2.6x 2.5x 2.5x 2.2x 2.6x 2.8x
Inference Speedup 3.8x 2.4x 2.9x 4.7x 2.3x 2.6x 2.5x 2.5x 2.3x 2.7x 2.9x
Input Seq Length 64 64 128 128 128 256 128 128 64 512 160
a Matthew’s Correlation is reported for CoLA, F1-score for MRPC and QQP, Pearson Correlation for STS-B, and Accuracy for the remaining tasks.
b Using equation (7).
Table III compares the proposed approach and the previously reported speedup methods. BERT-base model ($L$=12, $h$=12, $d$=768) is used as the basis for comparison. The test results, the average accuracy drop compared to BERT-base, the average number of FLOPs, and the average inference speedup are reported for each model. As shown in the table, the proposed approach achieves 2.9 times speedup with 0.73% on average accuracy drops compared to BERT-base model.
TABLE III: Comparison of Results Between BERTbase [3], DistilBERT [49], TinyBERT [50], MobileBERT [51], PoWER-BERT [28],
LayerDrop [42], FPT [37], and Latency-Adjustable BERTbase (This Work)
Modela Method CoLA SST-2 MRPC QQP QNLI RTE MNLIM MNLIMM Avg Drop FLOPs Speedup
BERT-base - 52.7 93.5 88.7 71.2 91.0 68.9 84.7 83.8 0.00 22.5B 1.0x
DistilBERT KD 51.3 91.3 87.5 70.1 89.2 59.9 82.2 - 2.74 11.3B 1.7x
TinyBERT KD 51.1 93.1 87.3 71.6 90.4 70 84.6 83.2 0.40 11.3B 2.0x
MobileBERTb KD 50.5 92.8 88.8 70.2 90.6 66.2 83.3 82.6 1.19 5.7B 5.5x
PoWER-BERTc Pruning 52.3 92.1 88.1 70.2 90.1 67.4 83.8 83.1 0.92 - 2.9x
LayerDrop Pruning - 92.5 - - 89.4 - 82.9 - 1.47 11.3 1.7x
FPT Pruning - 92.6 - - 90.3 - 83.5 - 0.93 14.5 1.5x
This Work Pruning 53.4 92.6 88.0 70.1 90.2 67.9 83.7 82.7 0.73 7.7B 2.9x
a Matthew’s Correlation is reported for CoLA, F1-score for MRPC and QQP, Pearson Correlation for STS-B, and Accuracy for the remaining tasks.
b MobileBERT starts from the BERT-large architecture, which has approximately three times as many parameters as BERT-base.
c PoWER-BERT training involves three fine-tuning steps, and the model’s inference speedup is fixed during inference.
@@
The table shows that the results of the suggested technique are competitive and more effective than previous baseline methods across a wide range of experimental tasks. Furthermore, compared to recently reported methods, our approach is applied only during the fine-tuning phase and offers the offline-tuning property. Blank fields in the table, indicate instances where some works did not report their assessments on certain datasets.
To assess the impact of the proposed method on GPT-2 model, the training process involves utilizing 1% of the OpenWebText [94] dataset with a stride of 128 for three epochs. Table IV exhibits the classification accuracy findings of the proposed method in contrast to GPT-2 and DistilGPT2. Notably, the proposed method attains an inference latency three times quicker than baseline model, accompanied by a marginal average accuracy reduction of 0.55%. Blank fields in the tables, indicate instances where some works did not report their assessments on certain datasets.
TABLE IV: Accuracy comparison between GPT-2 [73] and DistilGPT2 [49] with latency adjustable GPT-2 (This Work)
Model SST-2 IMDB Avg Drop Speedup
GPT-2 93.2 94.1 0.00 1.0x
DistilGPT2 91.5 92.1 1.85 1.9x
This Work 92.8 93.4 0.55 3.0x
IV-E Experimental Results on Text Generation Tasks
Table V presents the results of the zero-shot perplexity assessment for the suggested method in text generation tasks, offering a basis for comparison with other contemporary studies. The findings demonstrate that the method effectively maintains the global context of the input sequence. The proposed Attention Context Contribution (ACC) metric performs well, ensuring that the elimination of word-vectors does not compromise the input’s global context. It is noteworthy that the original GPT-2 model, trained by OpenAI, benefited from more extensive hardware and dataset resources. However, our proposed architecture enables efficient adaptation with just a few small epochs. Furthermore, the proposed technique exhibits superior performance when compared to the DistilGPT2 [49] and ZipGPT2 [95] models in the evaluation.
TABLE V: Perplexity (PPL) Comparison Between GPT-2 [73], DistilGPT2 [49], ZipGPT2 [95], and Latency-Adjustable GPT-2